OpenStudy (anonymous):

find the area of the rectangle The width is 11 length is 4 and 1over2

OpenStudy (anonymous):

Show work

OpenStudy (anonymous):

11*9/2=99/2=49.5

OpenStudy (anonymous):

first step is to turn mixed number into an improper faction. do you know how to do this?

OpenStudy (anonymous):

\[4\tfrac{1}{2}=\frac{2\times 8+1}{2}=\frac{9}{2}\]

OpenStudy (anonymous):

then multiply, get \[11\times \frac{9}{2}=\frac{99}{2}=49.5\] what dipanker said